### v1

Abstract
This paper studies the axiomatic bargaining problem and proposes a new class of bargaining solutions, called coarse Nash solutions. These solutions assign to each problem a set of outcomes coarser than that chosen by the classical Nash solution (nash1950bargaining). Our main result shows that these solutions can be characterized by new rationality axioms for choice correspondences, which are modifications of Nash’s independence of irrelevant alternatives (or more precisely, arrow1959rational’s (arrow1959rational) choice axiom), when combined with standard axioms.
Keywords: Axiomatic bargaining, Nash solution, Independence of irrelevant alternatives
JEL Classifications: D31, D63, D74
1 Introduction
Consider a group of individuals facing a collective choice problem, in which each individual only cares about the payoff of the chosen outcomes. In a seminal paper, nash1950bargaining formalized these situations as bargaining problems (sets of feasible utility vectors) and studied mappings $F$ that indicate the chosen outcomes, called bargaining solutions. Nash introduced several desirable axioms for bargaining solutions and showed that the Nash solution, which selects alternatives that maximize the product of the individuals’ utility levels of the outcomes for each feasible set, is the unique solution that satisfies all of them.
A key axiom for the Nash solution is arrow1959rational’s choice axiom, a set-valued version of Nash’s independence of irrelevant alternatives (cf. kaneko1980extension; xu2006alternative). This axiom states that if the feasible set expands and the rule still chooses some alternatives from the original set, then all originally chosen alternatives must still be selected (i.e., for two problems $S,S^{\prime}$ such that $S\subseteq S^{\prime}$, $F(S^{\prime})\cap S\neq\emptyset$ implies $F(S)=F(S^{\prime})\cap S$). This property is closely related to the rationalizability of multi-valued bargaining solutions or choice correspondences via welfare functions or orderings (ok1999revealed).
Arrow’s axiom has been criticized from various perspectives, leading to the proposal of weaker rationality axioms for choice correspondences and alternative solution concepts. It is well-known that Arrow’s axiom can be decomposed into two weaker axioms: the Chernoff axiom and the dual Chernoff axiom. Suppose that a choice set $S$ expands to $S^{\prime}$ and the rule still selects some outcomes from $S$. The Chernoff axiom requires that if the chosen outcomes in $S^{\prime}$ are feasible in $S$, then they must also be chosen in $S$ (i.e., $F(S^{\prime})\cap S\neq\emptyset$ implies $F(S)\supseteq F(S^{\prime})\cap S$). The dual Chernoff axiom requires that whenever some outcome chosen in $S$ is selected in $S^{\prime}$, all outcomes chosen in $S$ must also be done so in $S^{\prime}$ (i.e., $F(S^{\prime})\cap S\neq\emptyset$ implies $F(S)\subseteq F(S^{\prime})\cap S$). Most criticisms of Arrow’s axiom target the Chernoff axiom. In contrast, we argue that the dual Chernoff axiom is also subject to critique. [^1] For instance, Thomson1994CHAP questioned Nash’s independence of irrelevant alternatives under the name of contraction independence as follows: “(I)f the contraction described in the hypothesis of contraction independence is skewed against a particular agent, why should the compromise be prevented from moving against him?” This criticism is related to whether the outcome chosen in the original problem should remain chosen in a smaller problem, and this relates to the Chernoff axiom. Similar arguments have been made by many other studies, such as roth1979axiomatic and luce2012games. Motivated by these criticisms, various alternative solution concepts, such as the Kalai-Smorodinsky solution (kalai1975other), have been introduced.
To illustrate, suppose that $a$ and $b$ are available in $S$ and that both are chosen because they are incomparable. If $c$ becomes available in the expanded problem $S^{\prime}=S\cup\{c\}$ and $c$ is preferred to $a$ but incomparable to $b$, then choosing $b$ and $c$ in $S^{\prime}$ but not $a$ seems reasonable. However, the dual Chernoff axiom rules out such a decision because it requires that if $b$ is chosen, then $a$ should also be chosen.
This condition is naturally justified when an explicit objective function exists, as in the Nash solution, because the chosen outcomes coincide with those that dominate the remaining outcomes in terms of the objective function. However, constructing a single objective function becomes difficult when multiple criteria must be considered. For example, difficulties in aggregating diverse value judgments, such as efficiency and equity, or attempting to determine the bargaining power of individuals in real-life negotiations often resist reduction to a single objective function. In such situations, the dominance relation itself may remain incomplete, and the solution concept should allow for greater flexibility or a certain “buffer.”
This observation raises a natural question: What rationality axioms should we impose to construct collective choice rules, and what class of rules satisfies them? To address this issue, we propose two axioms by weakening Arrow’s choice axiom. The first axiom modifies the dual Chernoff axiom while retaining the Chernoff axiom. Recall that the original dual Chernoff axiom implies that if an outcome is chosen in both $S$ and $T$, it should also be chosen in $S\cup T$. Our axiom also requires the converse: If an outcome is chosen in $S\cup T$ and is feasible in both $S$ and $T$, it must be chosen in both. We call this axiom the weak Arrow axiom because it coincides with the combination of the Chernoff axiom and the weak dual Chernoff axiom known in the literature. The second axiom postulates that all outcomes chosen in the original problem should be selected in an expanded problem if no outcome that becomes feasible owing to the expansion is chosen. This prevents the disappearance of all chosen outcomes upon irrelevant expansions; we therefore call this axiom independence of irrelevant expansions. We investigate implications of these axioms under standard axioms such as Pareto optimality, scale invariance, and continuity.
Our main result shows that, under these axioms, the dominance relation induced by a solution can be generalized. Recall that the dominance relation in the Nash solution is represented by comparisons of the products of individuals’ utility levels. Equivalently, $x$ dominates $y$ if the log-transformed difference in their utility vectors lies in the set $\{z\in\mathbb{R}^{n}\mid\sum_{i=1}^{n}z_{i}>0\}$. Our representation theorem generalizes this set to any open set $A$, which we call an improving set, satisfying two conditions: (i) $A$ contains all strictly positive vectors and no weakly negative vectors (monotonicity), and (ii) for all $x,y\in A$, $x+y\in A$ (two-stage improvement property).
This class of solutions includes several meaningful examples. At one extreme are the weighted Nash solutions, while at the other extreme is the solution that assigns all weakly Pareto optimal outcomes. If $A$ is a convex cone, the dominance relation corresponds to the unanimity rule over multiple weighted Nash solutions. Since these solutions in the main result are less decisive than the original Nash solution but still based on the comparison of the Nash product, we call them coarse Nash solutions.
Furthermore, we formally show that these solutions yield larger choice sets than weighted Nash solutions: More precisely, for every coarse Nash solution, there exists a weight vector such that its chosen set always includes the outcomes chosen by the corresponding weighted Nash solution. Thus, each coarse Nash solution is literally coarser than some weighted Nash solution.
This paper is organized as follows. Section 2 formalizes the bargaining problem and the basic axioms satisfied by the Nash solution. Section 3 introduces axioms motivated by the limitations of Arrow’s choice axiom. Section 4 examines the implications of these axioms under standard conditions and shows that any solution satisfying them can be represented as a coarse Nash solution. Section 5 discusses the related literature. Finally, Section 6 concludes the paper.
2 The bargaining problem
Let $N=\{1,2,\ldots,n\}$ be the fixed set of players, where $n\geq 2$. If they reach a unanimous agreement, then they attain a utility vector $x=(x_{1},x_{2},\ldots,x_{n})\in\mathbb{R}_{++}^{n}$, where $x_{i}$ is player $i$’s utility level. [^2] Let $\mathbb{R}$ (resp. $\mathbb{R}_{+}$, $\mathbb{R}_{++},$ and $\mathbb{R}_{-}$) denote the set of real numbers (resp. nonnegative numbers, positive numbers, and nonpositive numbers). Let $\mathbb{R}^{n}$ (resp. $\mathbb{R}^{n}_{+}$, $\mathbb{R}^{n}_{++}$, and $\mathbb{R}^{n}_{-}$) denote the $n$-fold Cartesian product of $\mathbb{R}$ (resp. $\mathbb{R}_{+}$, $\mathbb{R}_{++}$, and $\mathbb{R}_{-}$). Let $\mathbb{N}$ be the set of natural numbers and $\mathbb{Q}$ be the set of rational numbers. If they fail to reach an agreement, then they stay at the origin $\mathbf{0}=(0,0,\ldots,0)\in\mathbb{R}^{n}$. Here, we assume that all agreements give the players positive utility levels. A (bargaining) problem $S\subseteq\mathbb{R}^{n}_{++}$ is a set of utility vectors attainable through unanimous agreements. We assume that each problem $S$ is compact relative to $\mathbb{R}^{n}_{++}$ and comprehensive (i.e., for all $x,y\in\mathbb{R}^{n}_{++}$ with $x\geq y$, $x\in S$ implies $y\in S$). [^3] For $x,y\in\mathbb{R}^{n}$, we write $x\gg y$ if $x_{i}>y_{i}$ for all $i\in N$, and $x\geq y$ if $x_{i}\geq y_{i}$ for all $i\in N$. Let $\mathcal{B}$ be the set of problems. For any $A\subseteq\mathbb{R}_{++}$, let
$\text{cmp}A=\{y\in\mathbb{R}^{n}_{++}\mid\text{there exists}~x\in A~\text{such that}~x\geq y\}$
be the comprehensive hull of $A$, that is, the smallest comprehensive set including $A$ (within $\mathbb{R}_{++}$).
We introduce several notations for convenience. A permutation is a bijection on $N$. Let $\Pi$ be the set of permutations. For $x\in\mathbb{R}^{n}$ and $\pi\in\Pi$, define $x_{\pi}$ as the vector $(x_{\pi(1)}.x_{\pi(2)},\ldots,x_{\pi(n)})\in\mathbb{R}^{n}$. A set $S\subseteq\mathbb{R}^{n}$ is said to be symmetric if $S=\{x_{\pi}\mid x\in S\}$ for all $\pi\in\Pi$. For $a,x\in\mathbb{R}^{n}$, let $a\ast x=(a_{1}x_{1},a_{2}x_{2},\ldots,a_{n}x_{n})$. Similarly, for $a\in\mathbb{R}^{n}$ and $S\subseteq\mathbb{R}^{n}$, let $a\ast S=\{a\ast x\mid x\in S\}$.
A (bargaining) solution $F$ assigns a nonempty subset $F(S)$ of $S$ to each problem $S\in\mathcal{B}$. Let $\Delta_{N}\subseteq\mathbb{R}^{n}_{+}$ be the set of weights over the players. For $w\in\Delta_{N}$, the weighted Nash solution associated with $w$, denoted by $F^{w}$, is the solution such that for all $S\in\mathcal{B}$,
$F^{w}(S)=\mathop{\rm arg~max}\limits_{x\in S}\prod_{i\in N}x_{i}^{w_{i}}.$ (1)
A special case with equal weight (i.e., $w_{i}=w_{j}$ for all $i,j\in N$) is the Nash solution $F^{\text{N}}$. That is, for all $S\in\mathcal{B}$,
$F^{\text{N}}(S)=\mathop{\rm arg~max}\limits_{x\in S}\prod_{i\in N}x_{i}.$ (2)
In other words, the Nash solution chooses the maximizers of the product of the player’s utility levels for each problem. The following are standard axioms and their variants that the Nash solution satisfies in our setup.
Arrow Axiom.
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\cap S\neq\emptyset$, then $F(S)=F(S^{\prime})\cap S$.
Efficiency.
For all $S\in\mathcal{B}$ and $x\in F(S)$, there is no $y\in S$ such that $y\gg x$. Furthermore, for all $x^{\prime}\in S$, if $x^{\prime}\geq y^{\prime}$ for some $y^{\prime}\in F(S)$, then $x^{\prime}\in F(S)$.
Anonymity.
For all symmetric problems $S\in\mathcal{B}$ and $x\in S$, if $x\in F(S)$, then $x_{\pi}\in F(S)$ for all $\pi\in\Pi$.
Scale Invariance.
For all $S\in\mathcal{B}$, $a\in\mathbb{R}^{n}_{++}$, and $x\in S$, we have $x\in F(S)$ if and only if $a\ast x\in F(a\ast S)$.
Continuity.
For all $S\in\mathcal{B}$ and $x\in S$, if there exist $\{S^{k}\}_{k\in\mathbb{N}}\subseteq\mathcal{B}$ and $\{x^{k}\}_{k\in\mathbb{N}}\subseteq\mathbb{R}_{++}^{n}$ such that (i) $x^{k}\in F(S^{k})$ for all $k\in\mathbb{N}$; (ii) $\{S^{k}\}_{k\in\mathbb{N}}$ converges to $S$ in the Hausdorff topology; and (iii) $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $x$ in $\mathbb{R}_{++}^{n}$, then $x\in F(S)$.
Arrow axiom (arrow1959rational) is a set-valued version of Nash’s independence of irrelevant alternatives (IIA). Under the usual interpretation, this axiom requires that if a feasible set becomes smaller, then the set of chosen outcomes should equal the restriction of the originally chosen set. In efficiency, we require that any chosen outcome should not be strictly Pareto dominated, and that if an outcome is better than some chosen outcome for every player, then it should also be chosen. This ensures that at least one (strictly) Pareto efficient outcome is chosen. Interpretations of the remaining axioms are standard.
3 Rationalizabiliy
In the classical characterization of the Nash solution, IIA plays a crucial role and provokes enormous discussions about the Nash solution. For the set-valued version of the Nash solution, Arrow axiom is a natural counterpart and plays the same crucial role (kaneko1980extension; xu2006alternative). It is known that this axiom can be decomposed into the following two parts:
Chernoff Axiom.
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\cap S\neq\emptyset$, then $F(S)\supseteq F(S^{\prime})\cap S$.
Dual Chernoff Axiom.
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\cap S\neq\emptyset$, then $F(S)\subseteq F(S^{\prime})\cap S$.
Based on this decomposition, the requirements of Arrow axiom can be understood as follows. Suppose that a choice set $S$ expands to $S^{\prime}$ and the rule still chooses some outcomes from the original problem $S$ (that is, $F(S^{\prime})\cap S\neq\emptyset$). Then, the first condition requires that the choice from the expanded problem $S^{\prime}$ that was available in the original problem should only be taken from the original chosen outcomes (i.e., Chernoff Axiom); and the second condition requires that all the chosen outcomes in the original problem $S$ must remain chosen (i.e., dual Chernoff Axiom).
In the literature, the first requirement has been criticized, and several alternative axioms have been introduced to propose more appropriate solution concepts. Here, however, we focus on the second condition and argue that it is not free from criticisms, too.
To illustrate this, we consider the following situations. Suppose that $a$ and $b$ are available in a choice set $S$ and that both options are chosen because they are incomparable. Now, imagine that another option $c$ becomes available by expanding the choice set to the problem $S^{\prime}$, and $c$ is more appealing than $a$. We also assume that there is still no clear relationship between $b$ and $c$. Then, choosing $b$ and $c$ but not $a$ in $S$ would be a natural decision. However, dual Chernoff axiom does exclude such choices.
In a social decision context, this intuition requires us to agree with the following scenarios: Suppose that outcome $E$ is equitable but offers relatively low utility levels to individuals, whereas outcome $U$ is unequal but provides a sufficiently high total utility. One problem with taking $U$ is that the redistribution of utility among individuals is not easily achieved due to political or ideological reasons. Now, suppose that a more appealing option $E^{\prime}$ than $E$ becomes feasible. For example, $E^{\prime}$ is an outcome in which the utility of every individual is higher than that under $E$, though its total utility is still lower than that under $U$. In such a situation, our common intuition suggests that $U$ and $E^{\prime}$ should be selected from the expanded choice set and $E$ should no longer be chosen. However, if we accept Arrow axiom, particularly dual Chernoff axiom, such a decision must be ruled out.
To avoid these counterintuitive conclusions stemming from dual Chernoff axiom, we consider the following weaker requirements:
Weak Dual Chernoff Axiom.
For all $S,T\in\mathcal{B}$, $F(S)\cap F(T)\subseteq F(S\cup T)$. [^4] This axiom is sometimes written as follows: For any $\{S^{k}\}_{k=1}^{K}\subseteq\mathcal{B}$, $\bigcap_{k=1}F(S^{k})\subseteq F(\bigcup_{k=1}^{K}S^{k})$. By induction, the equivalence of both axioms can be shown, so that we can use the present axiom without loss of generality.
Independence of Irrelevant Expansions (IIE).
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\subseteq S$, then $F(S)\subseteq F(S^{\prime})\cap S$.
Weak dual Chernoff axiom is already known in the literature on social choice theory. It requires that, for any problem $S$ and $T$, the commonly chosen outcomes in both problems should also be chosen in the combined problem $S\cup T$. This axiom is indeed weaker than dual Chernoff axiom because the latter implies both $F(S)\subseteq F(S\cup T)$ and $F(T)\subseteq F(S\cup T)$. Note that this axiom can accommodate the intuitive decision in the above scenarios. IIE is a new axiom about problem expansions. It requires that if the expansion is irrelevant in the sense that all the outcomes chosen in the expanded problem $S^{\prime}$ can be taken in the original problem $S$, then the outcomes chosen in $S$ should be chosen in $S^{\prime}$ again. This axiom avoids the above criticism of Arrow axiom by strengthening the prerequisite of dual Chernoff axiom.
As weaker rationality axioms for bargaining solutions than Arrow axiom, we adopt Chernoff axiom, weak dual Chernoff axiom, and IIE, and examine their implications under the aforementioned standard axioms. We first show that the implications of the combination of Chernoff axiom and weak dual Chernoff axiom are summarized as follows:
Proposition 1.
A bargaining solution satisfies Chernoff axiom and weak dual Chernoff Axiom if and only if it satisfies the following property:
For all $S,T\in\mathcal{B}$,   $F(S)\cap F(T)=F(S\cup T)\cap S\cap T$.
[…]
Proposition 2.
Suppose that a solution $F$ satisfies efficiency and continuity. Then, the following two statements hold.
  • (1)
[…]
    $F(S)=\{x\in S\mid x\succsim y,\forall y\in S\},~~\forall S\in\mathcal{B}.$
  • (2)
[…]
    $F(S)=\{x\in S\mid\not\exists y\in S~~\text{s.t.}~~y\succ x\},~~\forall S\in\mathcal{B}.$
Remark. In the finite-alternative case, sen1971choice shows the counterpart result of (2) in Proposition 2, using Chernoff axiom, weak dual Chernoff axiom, and so-called Sen’s $\delta$. [^5] In our model, it can be said that $F$ satisfies Sen’s $\delta$ if, for all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $\{x,y\}\subseteq F(S)$, then $\{x\}\neq F(S^{\prime})$. A stronger version of the axiom can be stated that, for all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $A\subseteq F(S)$, then $B\neq F(S^{\prime})$ for all $B\subsetneq A$. Even under this stronger version, our result cannot be established. Hence, technically speaking, this result is a generalization of Sen’s result to an infinite-alternative case by using a different axiom–IIE–with efficiency and continuity to address the additional complexity caused by infinitely many alternatives.
4 The coarse Nash solutions
This section examines the implications of the axioms introduced in the previous section by providing a representation theorem for them.
[…]
$x\succsim^{*}y\iff(e^{x_{1}},\ldots,e^{x_{n}})\succsim(e^{y_{1}},\ldots,e^{y_{n}}).$
In other words, for all $x,y\in\mathbb{R}^{n}_{++}$, we have
$\log x\succsim^{*}\log y\iff x\succsim y,$
[…]
$\log x\succsim^{*}\log y\iff(\log x-\log y)\succsim^{*}\textbf{0}\iff\sum_{i\in N}w_{i}(\log x_{i}-\log y_{i})\geq 0.$
That is, for any $S\in\mathcal{B}$, $x\in F(S)$ if and only if there is no $y\in S$ such that
$\log y-\log x\in\quantity{z\in\mathbb{R}^{n}\mid\sum_{i\in N}w_{i}z_{i}\geq 0}$ (3)
(Figure 1). Note that this is an alternative representation of the weighted Nash solution $F^{w}$, as we can check immediately. Here, if we regard a solution as an arbitrator’s choice rule, then (3) can be interpreted as follows: The benefit of the change from $x$ to $y$ is measured by the difference between their log-transformed vectors, and this change is regarded as a social improvement only if the difference is in the set $\{z\in\mathbb{R}^{n}\mid\sum_{i\in N}w_{i}z_{i}\geq 0\}$. The weighted Nash solution $F^{w}$ assigns to each problem the set of vectors that cannot be improved in this sense. If we consider a solution as the prediction of bargaining outcomes, whether the difference in the log-transformed vectors is in the set $\{z\in\mathbb{R}^{n}\mid\sum_{i\in N}w_{i}z_{i}\geq 0\}$ can be a criterion for determining which vector is more appropriate as a prediction.
[…]
Figure 1: The set $\left\{z\in\mathbb{R}_{++}^{n}\mid\sum_{i\in N}w_{i}z_{i}>0\right\}$ for some $w\in\Delta_{N}$
Conceptually, the right-hand side of (3) need not be the set $\{z\in\mathbb{R}^{n}\mid\sum_{i\in N}w_{i}z_{i}\geq 0\}$; we can take other sets to determine whether a social improvement is achieved or not. Our main result shows that if we replace Arrow axiom with weak Arrow axiom and IIE, then this set can be generalized, and the Nash solution can be modified to accommodate choice with some “buffer.”
[…]
Definition 1.
For an improving set $A\subseteq\mathbb{R}^{n}$, a solution is a coarse Nash solution associated with $A$, denoted by $F^{A}$, if for all $S\in\mathcal{B}$,
$F^{A}(S)=\{x\in S\mid\nexists y\in S~~\text{s.t. }\log y-\log x\in A\}.$ (4)
In the definition of coarse Nash solutions, the condition (i) of improving sets corresponds to the monotonicity of the dominance relation, while the condition (ii) means that two-stage improvements are also deemed as desirable changes.
We then present our main result. The following theorem shows that the coarse Nash solutions characterize all the implications of weak Arrow axiom, IIE, and other basic axioms.
Theorem 1.
A solution $F$ satisfies weak Arrow axiom, IIE, efficiency, scale invariance, and continuity if and only if it is a coarse Nash solution.
If anonymity is additionally imposed, the improving set becomes symmetric, as stated in the following corollary:
Corollary 1.
A solution $F$ satisfies weak Arrow axiom, IIE, efficiency, scale invariance, anonymity, and continuity if and only if it is a coarse Nash solution associated with a symmetric improving set.
To understand what solutions are included in the class of coarse Nash solutions, we provide several examples (and counterexamples) of them.
[…]
(a) The set $A$ in Example 3
[…]
(b) The set $A$ in Example 4
Figure 2: Examples of improving sets
Example 1.
Suppose that $A=\{z\in\mathbb{R}^{n}\mid\sum_{i\in N}w_{i}z_{i}>0\}$ for some $w\in\Delta_{N}$. As discussed earlier, the coarse Nash solution $F^{A}$ coincides with the weighted Nash solution $F^{w}$, which is an extreme case.
Example 2.
For the other extreme case, suppose that $A$ is the smallest set, that is $A=\mathbb{R}_{++}^{n}$. Then, for all $x,y\in\mathbb{R}_{++}^{n}$,
$\log y-\log x\in A$ $\iff[~\log y_{i}-\log x_{i}>0~~\text{for all }i\in N~]\iff[~y_{i}>x_{i}~~\text{for all }i\in N~],$
which implies that $F^{A}(S)=\{x\in S\mid\nexists y\in S~~\text{s.t. }y\gg x\}$ for all $S\in\mathcal{B}$. That is, $F$ is the solution that assigns to each problem the set of all weakly Pareto optimal agreements.
[…]
(a) The set $A^{\prime}$ in Example 3
[…]
(b) The set $A^{\prime}$ in Example 4
Figure 3: Counterexamples of improving sets
Example 3.
Suppose that an improving set $A$ is the convex cone $\{x\in\mathbb{R}^{n}_{++}\mid\sum_{i\in N}w_{i}x_{i}>0\quad(\forall w\in W)\}$ for some nonempty closed subset $W\subseteq\Delta_{N}$ (Figure 2(a)). In this case, for all $x,y\in\mathbb{R}_{++}^{n}$,
$\log y-\log x\in A\iff\quantity[~\prod_{i\in N}y^{w_{i}}_{i}>\prod_{i\in N}x_{i}^{w_{i}}\quad\forall w\in W~],$ (5)
which implies that $F(S)=\{x\in S\mid\nexists y\in S~~\text{s.t. }\prod_{i\in N}y^{w_{i}}>\prod_{i\in N}x^{w_{i}}_{i}~~(\forall w\in W)\}$ for all $S\in\mathcal{B}$. That is, the solution becomes a multi-weighted Nash solution, which chooses outcomes that can be justified by some set $W$ of weights.
By contrast, the set $A^{\prime}=\{x\in\mathbb{R}^{n}_{++}\mid\exists w\in W~~~\text{s.t.}~~~\sum_{i\in N}w_{i}x_{i}>0\}$ for some nonempty closed subset $W\subseteq\Delta_{N}$ with $|W|\geq 2$ is not an improving set (Figure 3(a)). For this set, we can take $x,y\in A^{\prime}$ such that $x+y\in\mathbb{R}^{n}_{-}$, which violates the definition of improving sets. Thus, the solution $F$ defined as $F(S)=\{x\in S\mid\nexists y\in S~~\text{s.t. }\prod_{i\in N}y^{w_{i}}>\prod_{i\in N}x^{w_{i}}_{i}~~(\exists w\in W)\}$ for all $S\in\mathcal{B}$ is not a coarse Nash solution.
The final example shows that there is no logical relationship between improving sets and the convexity.
Example 4.
Suppose $A=\mathbb{R}^{n}_{++}\cup\{x\in\mathbb{R}^{n}\mid\sum_{i\in N}x_{i}>\varepsilon\}$ for some $\varepsilon\in\mathbb{R}_{++}$ (Figure 2(b)). Note that this nonconvex set satisfies the definition of improving sets. Then, for all $x,y\in\mathbb{R}_{++}^{n}$,
$\log y-\log x\in A\iff[~y_{i}>x_{i}~~\text{for all }i\in N~]~\lor~\quantity[~\prod_{i\in N}y_{i}-\prod_{i\in N}x_{i}>\varepsilon~].$ (6)
Thus, $F^{A}$ can be represented as the solution that assigns to each problem $S\in\mathcal{B}$ the set of $x$ such that there is no $y\in S$ that satisfies (6). Note that the latter condition in (6) can be interpreted as the dominance relation associated with the Nash solution with a threshold $\varepsilon$: According to this, $y$ is deemed to be better than $x$ if the difference in their Nash products is greater than $\varepsilon$.
Furthermore, consider the convex set $A^{\prime}\subseteq\mathbb{R}^{n}$ drawn in Figure 3(b). Then, for some $\in A^{\prime}, x+x$$x\notin A^{\prime}$ holds, which violates the condition (ii) in the definition of improving sets.
From these examples, one might think that every coarse Nash solution yields an improving set that is sharper than that of some weighted Nash solution. The following result shows this conjecture is true and justifies that we call them coarse Nash solutions: For any coarse Nash solution, there exists a weighted Nash solution whose chosen outcomes are always included within those of the coarse solution.
Proposition 3.
For any improving set $A$, there exists $w\in\Delta_{N}$ such that $A\subseteq\{x\in\mathbb{R}^{m}\mid\sum_{i\in N}w_{i}x_{i}>0\}$. Furthermore, for all $S\in\mathcal{B}$, $F^{w}(S)\subseteq F^{A}(S)$ holds.
Proof.
Let $A$ be an arbitrary improving set. Since $A$ is an open set, it is sufficient to prove that for some $w^{*}\in\Delta_{N}$, $\textbf{co}(A)\subseteq\{x\in\mathbb{R}^{m}\mid\sum_{i\in N}w^{*}_{i}x_{i}\geq 0\}$ holds, where $\textbf{co}(A)$ is the convex hull of $A$. If $\mathbf{0}$ is not an interior point of $\textbf{co}(A)$, then by the separating hyperplane theorem, there exists $w^{*}\in\mathbb{R}^{n}$ such that $\sum_{i\in N}w^{*}_{i}x_{i}\geq 0$ for all $x\in\textbf{co}(A)$. Since $\mathbb{R}^{n}_{+}\subseteq A$, we can set $w^{*}\in\Delta_{N}$.
To complete the proof, we show that $\mathbf{0}$ is not an interior point of $\textbf{co}(A)$. Suppose to the contrary that $\mathbf{0}$ is an interior point of $\textbf{co}(A)$. Then, there exist $x^{1},x^{2},\ldots,x^{K}\in A$ and $\alpha^{1},\alpha^{2},\ldots,\alpha^{K}\in[0,1]$ with $\sum_{k=1}^{K}\alpha^{k}=1$ such that
$\sum_{k=1}^{K}\alpha^{k}x^{k}\in\mathbb{R}^{n}_{-}.$
Note that we can take $\alpha^{1},\alpha^{2},\ldots,\alpha^{K}\in\mathbb{Q}\cap[0,1]$, that is, there exist $p^{1},p^{2},\ldots,p^{K},q^{1},q^{2},\ldots,q^{K}\in\mathbb{N}$ such that $\alpha^{k}=p^{k}/q^{k}$ for each $k=1,2,\ldots,K$. Let $x^{*}=(\prod_{k=1}^{K}q_{k})\sum_{k=1}^{K}p^{k}x^{k}/q^{k}\in\mathbb{R}^{n}_{-}$. By applying the property (ii) in the definition of improving sets repeatedly, we have $x^{*}\in A$, which is a contradiction to the property (i) in the definition of improving sets. ∎
Putting differently, this proposition states that for each coarse Nash solution $F^{A}$, there exists $w\in\Delta_{N}$ such that for all $x,y\in\mathbb{R}^{n}_{++}$,
$x\succ y\implies\prod_{i\in N}x_{i}^{w_{i}}\geq\prod_{i\in N}y_{i}^{w_{i}},$ (7)
[…]
Furthermore, from this result, we can obtain the following important properties of the collection of improving sets.
Corollary 2.
For the collection of improving sets, $\mathbb{R}^{n}_{++}$ is the minimum set with respect to the set inclusion, and a set $A^{*}$ is a maximal set with respect to the set inclusion if and only if $A^{*}=\{x\in\mathbb{R}^{m}\mid\sum_{i\in N}w_{i}x_{i}>0\}$ for some $w\in\Delta_{N}$.
If we additionally require that an improving set $A$ be symmetric, a weight vector specified in Proposition 3 can be taken to be symmetric. That is, under the symmetry of $A$, the solution $F^{A}$ assigns a coarser set of outcomes than that chosen by the Nash solution.
Proposition 4.
If $A$ is a symmetric improving set, then $A\subseteq\{x\in\mathbb{R}^{n}\mid\sum_{i\in N}x_{i}>0\}$. Furthermore, for all $S\in\mathcal{B}$, $F^{\text{N}}(S)\subseteq F^{A}(S)$ holds.
Proof.
Let $A$ be a symmetric improving set. Suppose to the contrary that there exists $x\in A$ with $\sum_{i\in N}x_{i}\leq 0$. Then, by the symmetry of $A$, for each $\pi\in\Pi$, we have $x^{\pi}\in A$. Let $x^{\ast}=\sum_{\pi\in\Pi}x^{\pi}$, that is, $x^{\ast}=\bigl((n-1)!(\sum_{i\in N}x_{i}),\ldots,(n-1)!(\sum_{i\in N}x_{i})\bigr)$. By $\sum_{i\in N}x_{i}\leq 0$, we have $x^{\ast}\leq\mathbf{0}$. By applying the property (ii) in the definition of improving sets repeatedly, we have $x^{\ast}\in A$, which is a contradiction. ∎
5 Related literature
Motivated by the problems of Arrow axiom and Nash’s IIA, many alternative axioms and solution concepts, such as kalai1975other’s (kalai1975other) solution, have been proposed (for comprehensive surveys, see thomson1981nash; Thomson1994CHAP; Thomson2022RED). Nash’s IIA has been criticized for the property that the solution outcomes in the original problem must also be chosen in a smaller problem, even if the contraction dramatically changes the bargaining situations (e.g., luce2012games and roth1979axiomatic). Under the decomposition of Arrow axiom, Chernoff axiom has therefore attracted considerable attention. In contrast, this paper focuses on a problem that arises from dual Chernoff axiom. By weakening this part, we have obtained a class of generalized Nash solutions called coarse Nash solutions.
Relationships between the rationalizability of choice correspondences and choice axioms have been examined extensively in social choice theory. In a series of works, sen1970collective; sen1971choice; sen1977social examined the rationalizability of choice correspondences in an abstract setting. [^7] See cato2018choice for a detailed analysis of the relationships among several choice axioms. In the context of the bargaining problem, peters1991independence directly investigated the implications of Nash’s IIA in single-valued bargaining solutions. ok1999revealed showed that Arrow axiom plays a crucial role in the rationalizability of multi-valued bargaining solutions. nakamura2025wp; nakamura2025wp2 also considered a weakening of Arrow axiom, but adopted a different approach. Nakamura’s axiom postulates the requirement of Arrow axiom only when each player’s maximum utility level coincides in both the original and new problems (for similar axioms, see also DUBRA2001131 and xu2006alternative). Combined with standard axioms, especially scale invariance, this weakened choice axiom implies rationalizability under the 0–1 normalization of each player’s utility levels.
It should be noted that we employ a domain different from nash1950bargaining’s (nash1950bargaining) original formalization. In the most standard framework of the bargaining problem, feasible sets are assumed to be convex, and bargaining solutions are formalized as choice functions. By contrast, we do not assume the convexity of feasible sets and allow solutions to select multiple outcomes. Non-convex bargaining problems have been considered in several studies, such as kaneko1980extension, herrero1989nash, conley1991bargaining, zhou1997nash, ok1999revealed, mariotti2000maximal, xu2006alternative; xu2013rationality; xu2019equitable, and peters2012wpo. For surveys, see also xu2020nonconvex and Thomson2022RED. The convexity has been justified through the assumptions that randomization among agreements is allowed and that each player follows the expected utility theory. In the non-convex problem, we can analyze the situations in which these assumptions do not hold. Note that in such cases, scale invariance still has compelling interpretations, even though this axiom has traditionally been justified within the expected utility framework. For a detailed discussion, see xu2006alternative.
A coarse Nash solution can be regarded as the combination of maximizers of multiple choice criteria, as illustrated in the examples we discussed earlier. The idea of conjoining multiple choice criteria is closely related to tadenuma2002efficiency; tadenuma2005egalitarian and houy2009lexicographic, where the choice correspondence is defined through the sequential application of efficiency and fairness criteria in the context of general resource allocation problems. See also manzini2012choice for a general model of the sequential composition of different choice criteria. In contrast to this sequential approach, our choice correspondence takes into account the intersection of multiple criteria. Such intersections typically involve trade-offs, which often result in incompleteness and render the resulting predictions coarse (e.g., kamaga2018utilitarianism).
Coarse Nash solutions are also related to incomplete preferences studied in decision theory, such as the expected multi-utility model (dubra2004expected) and the multi-prior expected utility model (bewley2002knightian). These preferences can be represented by the unanimity rule among multiple parameters in the decision maker’s mind, and are characterized by a violation of completeness. The special cases considered in Example 3 correspond to these models. The class of coarse Nash solutions encompasses a wider range of rules, such as the solution in Example 4. This difference arises from the properties of improving sets, which are not necessarily convex, whereas the upper contour sets of the above decision-making models are convex.
Furthermore, Proposition 3 is related to aumann1962utility’s (aumann1962utility) result. Aumann showed that for any risk preference that satisfies the axioms of the expected utility theory except for completeness, there exists a von Neumann-Morgenstern utility function compatible with the preference in a weak sense. Proposition 3 shows that a similar result holds for a binary relation induced by any improving set: There exists a weight over players that is compatible with the binary relation. Despite their similarities, our proposition is not a direct application of Aumann’s result or argument because of the differences in the geometric properties of the upper contour sets: Unlike Aumann’s argument, improving sets are not necessarily convex cones, and hence the separating hyperplane theorem cannot cover all the cases in our analysis.
6 Concluding comments
This paper has focused on the limitations of Arrow axiom in the classical bargaining problem, particularly targeting its component, dual Chernoff axiom. We have introduced two new rationality axioms that weaken Arrow axiom: weak Arrow axiom and IIE. The main contribution of this paper is to show that these new axioms, when combined with standard axioms, characterize a new class of solutions called coarse Nash solutions.
Coarse Nash solutions generalize the dominance relation of the Nash solution by employing more general open sets, which we call improving sets. This class of solutions encompasses a wide range of outcomes, including weighted Nash solutions and solutions that assign all weakly Pareto optimal outcomes as extreme cases. As its name suggests, this solution always selects a larger (coarser) set of outcomes than some weighted Nash solution. Consequently, it provides a more flexible guideline for collective decision-making in real-world bargaining situations where multiple criteria, such as efficiency and equity, may conflict, making it difficult to construct a single objective function.
Appendix
A A Proof of Proposition 1
Suppose that $F$ satisfies Chernoff axiom and weak dual Chernoff axiom. Take any $S,T\in\mathcal{B}$. By weak dual Chernoff axiom, $F(S)\cap F(T)\subseteq F(S\cup T)$. Clearly, $F(S)\cap F(T)\subseteq S\cap T$. Therefore, $F(S)\cap F(T)\subseteq F(S\cup T)\cap S\cap T$ holds. Then, suppose that $F(S\cup T)\cap S\cap T\neq\emptyset$. Since $F(S\cup T)\cap S\neq\emptyset$ and $F(S\cup T)\cap T\neq\emptyset$, by Chernoff axiom, we have $F(S\cup T)\cap S\subseteq F(S)$ and $F(S\cup T)\cap T\subseteq F(T)$. Therefore, $F(S)\cap F(T)\supseteq F(S\cup T)\cap S\cap T$.
Conversely, suppose that $F$ satisfies the property of the proposition, weak Arrow axiom. Then, we show that $F$ also satisfies Chernoff axiom and weak dual Chernoff axiom. For Chernoff axiom, take $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$. By weak Arrow axiom, we have $F(S)\cap F(S^{\prime})=F(S^{\prime})\cap S$. By $F(S)\cap F(S^{\prime})\subseteq F(S)$, we have $F(S)\supseteq F(S^{\prime})\cap S$. For weak dual Chernoff axiom, take $S,T\in\mathcal{B}$. Then, by weak Arrow axiom, we have $F(S)\cap F(T)=F(S\cup T)\cap S\cap T\subseteq F(S\cup T)$. ∎
A B Proof of Proposition 2
Since the if part is clear in each point, we show the only-if part. Given a bargaining solution $F$, define a binary relation $\succsim^{F}$ on $\mathbb{R}^{n}_{++}$ such that
$x\succsim^{F}y\iff x\in F(\text{cmp}\{x,y\})$
for any $x,y\in\mathbb{R}^{n}_{++}$.
Step 1: If $F$ satisfies efficiency, then $\succsim^{F}$ is complete.
Take any $x,y\in\mathbb{R}^{n}_{++}$. Since $F$ is nonempty for any $S\in\mathcal{B}$, there exists $z\in\text{cmp}\{x,y\}$ such that $z\in F(\text{cmp}\{x,y\})$. By definition, $x\geq z$ or $y\geq z$. Therefore, by efficiency, we have $x\in F(\text{cmp}\{x,y\})$ or $y\in F(\text{cmp}\{x,y\})$, that is, $x\succsim^{F}y$ or $y\succsim^{F}x$ holds.
Step 2: If $F$ satisfies efficiency and continuity, then $\succsim^{F}$ is monotone and continuous.
The monotonicity follows from efficiency. To show the continuity, by Step 1, it suffices to prove that for all $x\in\mathbb{R}^{n}_{++}$, the set $U_{x}=\{y\in\mathbb{R}^{n}_{++}\mid y\succ^{F}x\}$ is open, or equivalently, $U_{x}^{c}=\{y\in\mathbb{R}^{n}_{++}\mid y\not\succ^{F}x\}=\{y\in\mathbb{R}^{n}_{++}\mid x\succsim^{F}y\}$ is closed. Let $x,z\in\mathbb{R}^{n}_{++}$ and $\{z^{k}\}_{k\in\mathbb{N}}\subseteq U_{x}^{c}$ be a sequence that converges to $z$. By the definition of $\succ$, $x\in F(\text{cmp}\{x,z^{k}\})$. By continuity, $x\in F(\text{cmp}\{x,z\})$, which means that $z\not\succ x$. Therefore, $U_{x}^{c}$ is a closed set.
Step 3: If $F$ satisfies weak Arrow axiom and IIE, then $\succsim^{F}$ is quasi-transitive.
Let $x,y,z\in\mathbb{R}^{n}_{++}$ be such that $x\succ^{F}y$ and $y\succ^{F}z$. By the definition of $\succ^{F}$, we have $x\in F(\text{cmp}\{x,y\})$, $y\notin F(\text{cmp}\{x,y\})$, $y\in F(\text{cmp}\{y,z\})$, and $z\notin F(\text{cmp}\{y,z\})$. We show that $x\in F(\text{cmp}\{x,y,z\})$, $y\notin F(\text{cmp}\{x,y,z\})$, and $z\notin F(\text{cmp}\{x,y,z\})$. Note that since $x\in F(\text{cmp}\{x,y,z\})$ follows from $y\notin F(\text{cmp}\{x,y,z\})$, $z\notin F(\text{cmp}\{x,y,z\})$, and efficiency, it is sufficient to prove the latter two ones. Suppose to the contrary that $y\in F(\text{cmp}\{x,y,z\})$ or $z\in F(\text{cmp}\{x,y,z\})$. If $y\in F(\text{cmp}\{x,y,z\})$, then by weak Arrow axiom, $y\in F(\text{cmp}\{x,y,z\})\cap\text{cmp}\{x,y\}\cap\text{cmp}\{x,y,z\}\subseteq F(\text{cmp}\{x,y\})$. This contradicts $y\notin F(\text{cmp}\{x,y\})$. On the other hand, if $z\in F(\text{cmp}\{x,y,z\})$, then by weak Arrow axiom, $z\in F(\text{cmp}\{x,y,z\})\cap\text{cmp}\{y,z\}\cap\text{cmp}\{x,y,z\}\subseteq F(\text{cmp}\{y,z\})$, which is also a contradiction with $z\notin F(\text{cmp}\{y,z\})$.
[…]
Step 4: If $F$ satisfies efficiency, continuity, weak Arrow axiom, and IIE, then it can be weakly rationalizable by $\succ^{F}$, that is,
$F(S)=\{x\in S\mid\not\exists y\in S~~\text{s.t.}~~y\succ^{F}x\},~~\forall S\in\mathcal{B}.$
Take $S\in\mathcal{B}$ arbitrarily. To prove $F(S)\subseteq\{x\in S\mid\nexists y\in S~~\text{s.t.}~~y\succ^{F}x\}$, suppose to the contrary that there exist $x,y\in S$ with $x\in F(S)$ but $y\succ x$. By weak Arrow axiom, $F(S\cup\text{cmp}\{x,y\})\cap S\cap\text{cmp}\{x,y\}=F(S)\cap F(\text{cmp}\{x,y\})$. Since $x$ is in the left hand side, $x\in F(\text{cmp}\{x,y\})$, a contradiction to $y\succ^{F}x$.
To prove the converse, let $x\in S$ be such that there is no $y\in S$ with $y\succ^{F}x$. By the definition of $\succsim^{F}$, $x\in F(\text{cmp}\{x,y\})$ or $y\notin F(\text{cmp}\{x,y\})$ for all $y\in S\setminus\{x\}$. However, if $y\notin F(\text{cmp}\{x,y\})$ for some $y\in S\setminus\{x\}$, by efficiency, we must have $x\in F(\text{cmp}\{x,y\})$. Hence,
$x\in F(\text{cmp}\{x,y\}),~~\forall y\in S.$ (8)
Let $\{y^{k}\}_{k\in\mathbb{N}}\subseteq S$ be a sequence such that $\{\text{cmp}\{x,y^{1},\ldots,y^{k}\}\}_{k\in\mathbb{N}}\subseteq\mathcal{B}$ converges to $S$. For all $K\in\mathbb{N}$, if $x\in F(\text{cmp}\{x,y^{1},\ldots,y^{K}\})$, then by weak Arrow axiom and (8), we have $x\in F(\text{cmp}\{x,y^{1},\ldots,y^{K}\})\cap F(\text{cmp}\{x,y^{K+1}\})\subseteq F(\text{cmp}\{x,y^{1},\ldots,y^{K},y^{K+1}\})$. This observation implies that for all $k\in\mathbb{N}$, $x\in F(\text{cmp}\{x,y^{1},\ldots,y^{k}\})$. By continuity, we have $x\in F(S)$.
Step 5: If $F$ satisfies efficiency, and Arrow axiom, then $\succsim^{F}$ is transitive.
Let $x,y,z\in\mathbb{R}^{n}_{++}$ be such that $x\succsim^{F}y$ and $y\succsim^{F}z$. By efficiency, $w\in F(\text{cmp}\{x,y,z\})$ for at least one $w\in\{x,y,z\}$. First, we show that $x\in F(\text{cmp}\{x,y,z\})$. Suppose that $y\in F(\text{cmp}\{x,y,z\})$. Then, by Arrow axiom, $F(\text{cmp}\{x,y\})=F(\text{cmp}\{x,y,z\})\cap\text{cmp}\{x,y\}$. Since $x\succsim^{F}y$, we have $x\in F(\text{cmp}\{x,y\})$, so $x\in F(\text{cmp}\{x,y,z\})$. Suppose that $z\in F(\text{cmp}\{x,y,z\})$. Similarly, by Arrow axiom, $F(\text{cmp}\{y,z\})=F(\text{cmp}\{x,y,z\})\cap\text{cmp}\{y,z\}$. Since $y\succsim^{F}z$, we have $y\in F(\text{cmp}\{y,z\})$, so $y\in F(\text{cmp}\{x,y,z\})$. Then, as the above argument, we can also see that $x\in F(\text{cmp}\{x,y,z\})$. Finally, by Arrow axiom, we can see that $x\in F(\text{cmp}\{x,y,z\})\cap\{x,z\}=F(\text{cmp}\{x,z\})$, that is, $x\succsim^{F}z$.
By Steps 1–5, we can obtain the result as follows. First, Steps 2, 3 and 4 directly imply that (2) of the proposition holds. Moreover, if $F$ satisfies Arrow axiom, by Steps 1 and 4,
$F(S)=\{x\in S\mid\not\exists y\in S~~\text{s.t.}~~y\succ^{F}x\}=\{x\in S\mid x\succsim^{F}y,\forall y\in S\},~~\forall S\in\mathcal{B}.$
Then, by Steps 1 and 5, $\succsim^{F}$ is complete and transitive, which is the assertion of (1). ∎
A C Proof of Theorem 1
We first show the only-if part. Let $F$ be a solution that satisfies weak Arrow axiom, IIE, efficiency, scale invariance, and continuity. By (2) of Proposition 2, we have known that
$F(S)=\{x\in S\mid\nexists y\in S~~\text{s.t.}~~y\succ x\},~~\forall S\in\mathcal{B},$
[…]
$x\succsim y\iff x\in F(\text{cmp}\{x,y\})$
for all $x,y\in\mathbb{R}^{n}_{++}$. Moreover, by scale invariance, it is easy to verify that $\succ$ is scale invariant (i.e., for all $a,x,y\in\mathbb{R}^{n}_{++}$, $x\succ y$ if and only if $a*x\succ a*y$). Furthermore, by the definition of $\succ$, it is asymmetric.
Next, we construct an improving set. For $x\in\mathbb{R}^{n}_{++}$, define $A_{x}\subseteq\mathbb{R}^{n}$ as
$A_{x}=\{(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\in\mathbb{R}^{n}\mid y\in\mathbb{R}_{++}^{n}~~\text{s.t.}~~y\succ x\}.$ (9)
We prove that for all $x,x^{\prime}\in\mathbb{R}^{n}_{++}$, $A_{x}=A_{x^{\prime}}$. For $z\in A_{x}$, there exists $y\in\mathbb{R}^{n}_{++}$ such that $y\succ x$ and $z=(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})$. Take $a\in\mathbb{R}^{n}_{++}$ as $a\ast x=x^{\prime}$. Since $\succ$ is scale invariant, $a\ast y\succ a\ast x=x^{\prime}$. By
$(\log a_{1}y_{1},\ldots,\log a_{n}y_{n})-(\log a_{1}x_{1},\ldots,\log a_{n}x_{n})$
$=(\log y_{1}+\log a_{1},\ldots,\log y_{n}+\log a_{n})-(\log x_{1}+\log a_{1},,\ldots,\log x_{n}+\log a_{n})$
$=(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})$
$=z,$
we have $z\in A_{x^{\prime}}$, which implies $A_{x}\subseteq A_{x^{\prime}}$. By a similar argument, we can prove the converse inclusion. Therefore, $A_{x}$ does not depend on $x$. Henceforth, we denote the set in (9) by $A$.
[…]
Finally, by the above discussion, we have shown that $y\succ x$ is equivalent to $(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\in A$ for any $x,y\in\mathbb{R}^{n}_{++}$ and $A$ is an improving set. Therefore, we can conclude that, for all $S\in\mathcal{B}$,
$F(S)$ $=$ $\{x\in S\mid\nexists y\in S~~\text{s.t.}~~y\succ x\}$
$=$ $\{x\in S\mid\nexists y\in S~~\text{s.t. }(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\in A\},$
which completes the proof of the only-if part.
Next, we prove the if part. It is easy to prove that the coarse Nash solutions satisfy efficiency. To verify the remained axioms, let $F^{A}$ be some coarse Nash solution.
  • Scale invariance: Let $S\in\mathcal{B}$, $a\in\mathbb{R}^{n}_{++}$, and $x\in F^{A}(S)$. By the definition of $F^{A}$, $F^{A}(S)=\{x\in S\mid\nexists y\in S~~\text{s.t. }(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\in A\}$. Therefore, by the definition of $a\ast S$,
    $F^{A}(a\ast S)$ $=\{x\in S\mid\nexists y\in S~~\text{s.t. }(\log a_{1}y_{1},\ldots,\log a_{n}y_{n})-(\log a_{1}x_{1},\ldots,\log a_{n}x_{n})\in A\}$
    $=\{x\in S\mid\nexists y\in S~~\text{s.t. }(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\in A\}$
    $=F^{A}(S).$
  • Continuity: Let $\{S^{k}\}_{k\in\mathbb{N}}\subseteq\mathcal{B}$ and $\{x^{k}\}_{k\in\mathbb{N}}\subseteq\mathbb{R}_{++}^{n}$ be such that (i) $x^{k}\in F^{A}(S^{k})$ for all $k\in\mathbb{N}$; (ii) $\{S^{k}\}_{k\in\mathbb{N}}$ converges to some $S\in\mathcal{B}$; (iii) $\{x^{k}\}_{k\in\mathbb{N}}$ converges to some $x\in S$ in $\mathbb{R}_{++}^{n}$. Take $z\in S$ arbitrarily. Let $\{z^{k}\}_{k\in\mathbb{N}}\subseteq\mathbb{R}^{n}_{++}$ be such that $z^{k}\in S^{k}$ for each $k\in\mathbb{N}$ and $\{z^{k}\}_{k\in\mathbb{N}}$ converges to $z$. By the definitions of $F^{A}$ and $\{x^{k}\}_{k\in\mathbb{N}}$, we have $(\log z^{k}_{1},\ldots,\log z^{k}_{n})-(\log x^{k}_{1},\ldots,\log x^{k}_{n})\notin A$ for each $k\in\mathbb{N}$. By the openness of $A$ and the continuity of the log function, $(\log z_{1},\ldots,\log z_{n})-(\log x_{1},\ldots,\log x_{n})\notin A$ holds. Since this holds for each $z\in A$, we have $x\in F^{A}(S)$.
  • Weak Arrow axiom: Let $S,T\in\mathcal{B}$ and $x\in\mathbb{R}^{n}_{++}$. If $x\in F^{A}(S)\cap F^{A}(T)$, then for all $y\in S\cup T$, $(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\not\in A$, which implies $x\in F^{A}(S\cup T)$. For the converse, suppose that $x\in F^{A}(S\cup T)\cap S\cap T$. Then, for all $y\in S\cup T$, $(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\not\in A$, which implies $x\in F^{A}(S)\cap F^{A}(T)$.
  • IIE: Let $S,S^{\prime}\in\mathcal{B}$ be such that $S\subseteq S^{\prime}$ and $F^{A}(S^{\prime})\subseteq S$. Take $x\in F^{A}(S)$ arbitrarily. By the definition of $F^{A}$,
    $(\log y_{1},\ldots,\log y_{n})-(\log x_{1},\ldots,\log x_{n})\not\in A,~~\forall y\in S.$ (10)
    If $x\notin F^{A}(S^{\prime})$, then there exists $z\in S^{\prime}\setminus S$ such that $(\log z_{1},\ldots,\log z_{n})-(\log x_{1},\ldots,\log x_{n})\in A$. We can set that $z$ is in $F^{A}(S^{\prime})$ by the property (ii) of improving sets and (10). This is a contradiction to $F^{A}(S^{\prime})\subseteq S$.